OpenStudy (anonymous):

Please help me again with this... (x+y)^2 ------- + y z if x = -1 y = 2 z = -2 z

OpenStudy (anonymous):

HAT?

OpenStudy (anonymous):

hello shinigami1m...it's an evaluation of algebra

OpenStudy (anonymous):

\[\frac{(x+y)^{2}}{z} + y = \frac{3}{2}\] if x = -1 y = 2 z = -2

OpenStudy (anonymous):

JohnnyRocket your answer is 3/2?

OpenStudy (anonymous):

\[\frac{-1+2}{-2}+2\]\[-\frac{1}{2}+2\]\[\frac{-1+4}{2}\]\[x=\frac{3}{2}\]

OpenStudy (anonymous):

Nancy Lam should I make it a mixed number because it is improper or 3/2 is okay when it comes to algebra?

OpenStudy (anonymous):

its just fine dear

OpenStudy (anonymous):

unless ur teacher told u to do so